A rectangle has width the same as side of a square whose perimeter is 20 m. the length of the rectangle is 9 m. Find the perimeter of this rectangle.
Answer by checkley71(8403)

WIDTH OF THE SQUARE=20/4=5M.
PERIMETER OF THE RECTANGLEIS:
2*5+2*9=10+18=28 FOR THE PERIMETER OF THE 5x9 RECTANGLE.